OpenStudy (anonymous):

x-2y+z=6 x+3y+z=16 3x-y-z=2 already solved: y=2

OpenStudy (anonymous):

i solved this but the server went down the roots are x = 3.5, y=2 and z = 6.5.

OpenStudy (anonymous):

this is what i got but somehow that's incorrect

OpenStudy (anonymous):

oh ive checked it out - lets try again

OpenStudy (anonymous):

no -those figures are correct for the equations you posted

OpenStudy (anonymous):

plus them into the equations

OpenStudy (anonymous):

alright thank u...must be a typo in the back of the book

OpenStudy (anonymous):

@jimmyrep, u answered my other question too but i was wondering if u can explain how u got the answer

jhonyy9 (jhonyy9):

x-2y+z=6 --- x-2y+z=6 x+3y+z=16 ---------------- x+3y+z=16 3x-y-z=2 ---- 3x-y-z=2 ---- 3x-y-z=2 ----------- -------------- 4x-3y=8 4x+2y=18 4x-3y=8 / x(-1) --- -4x+3y=-8 4x+2y=18 ---- 4x+2y=18 ------------ 0 5y=10 --- y=10/5 =2 if y=2 --- 4x-3(2)=8 --- 4x-6=8 --- 4x=14 --- x=14/4 =7/2 x-2y+z=6 --- z=6-x+2y=6-7/2 +4=10-7/2 2z=20-7 --- 2z=13 --- z=13/2 x=7/2 y=2 z=13/2